OpenStudy (tomfoolery1):

Using the range rule of thumb find the MAX and MIN usual values of the DATA set. x P(x) 0 0.005 1 0.016 2 0.042 3 0.116 4 0.198 5 0.245 6 0.202 7 0.112 8 0.038 9 0.011 10 0.015

OpenStudy (snowsurf):

Max and min are easy to find visually. So the best thing to do here is you can create a graph. You can find the max and min from just this table here too. Max will be the largest P(x) value and min will be the smallest P(x) value. So from this what are the min and max values?

OpenStudy (tomfoolery1):

My question seems to ask for the ST dev and mean of the data set and then using this equation mean +2 * the st dev to find the MAX value and mean - 2 * st dev to find the min value

OpenStudy (snowsurf):

You have a 11 data points in your set. To find the mean you add up all the data you have and divide by the total of data points you have. \[Mean = \frac{ \sum_{}^{} P(x) }{ N }\] Mean = 0.005 + 0.016 + 0.042 + 0.116 + 0.198 + 0.245 + 0.202 + 0.112 + 0.038 + 0.011 + 0.015 / 11 \[ST Dev = \sqrt{\frac{ 1 }{ N }\sum_{}^{}(x-mean)^{2}}\] Once you find the mean you then use the ST Dev formula. N is the number of data pints you have and x is your data point.